OpenStudy (anonymous):

Let f(x) = -4x + 7 and g(x) = 10x - 6. Find f(g(x)).

OpenStudy (kiakat):

since f(g(x)) you have to plug in the equation g(x) where ever x is in f(x)

OpenStudy (anonymous):

I got -40+31 is that correct?

OpenStudy (kiakat):

-40x+31, yes :)

OpenStudy (anonymous):

okay what about if you have an equation that has (f(g(2))?

OpenStudy (kiakat):

with the same equations for f(x) and g(x)??

OpenStudy (anonymous):

yes the problem is f(x)=x+8 and g(x)=x squared -6x-7

OpenStudy (kiakat):

okay for this one, the only thing you have to do differently, is plug in 2 wherever x is in g(x) first. then after that, you can plug in whatever the answer for g(2) is into wherever x is in f(x)

OpenStudy (kiakat):

the answer for this one wont have a variable.

OpenStudy (anonymous):

okay so I would make g(x) = 2 squared-12-7?

OpenStudy (kiakat):

yes.

OpenStudy (anonymous):

I got -15 but that is not an option

OpenStudy (kiakat):

thats not right. simplify g(x) before you put it into f(x)

OpenStudy (kiakat):

if g(2) that is. you get 2^2-12-7 or 4-12-7

OpenStudy (anonymous):

-15+8=-7?

OpenStudy (kiakat):

yep! is that one of the options??

OpenStudy (anonymous):

yes it is

OpenStudy (kiakat):

awesome. Do you think you could do those kinda of problems again without help now? I had trouble with those at first i think lol

OpenStudy (anonymous):

yeah thanks so much for your help

OpenStudy (kiakat):

No problem at all, anytime :)